OpenStudy (anonymous):

6m - 2 = 2m + 16 step by step:) thanks

hero (hero):

6m - 2 = 2m + 16 Subtract 2m from both sides to get: 6m - 2m - 2 = 16 Add 2 to both sides to get 6m - 2m = 16 + 2 Combine terms to get: 4m = 18 Divide both sides by 4 to get m = 18/4 Reduce the fraction by 2/2 to get m = 9/2

OpenStudy (anonymous):

Thanks:)